# The Streaming Batch Model for Efficient and Fault-Tolerant Heterogeneous Execution

### v4

@@ Abstract
While ML model training and inference are both GPU-intensive, CPU-based data processing is often the bottleneck. Distributed data processing systems based on the batch or stream processing models assume homogeneous resource requirements. They excel at CPU-based computation but either under-utilize heterogeneous resources or impose high overheads on failure and reconfiguration. We introduce the streaming batch model, a hybrid of the two models that enables efficient and fault-tolerant heterogeneous execution. The key idea is to execute one partition at a time to allow lineage-based recovery with dynamic resource allocation. This enables memory-efficient pipelining across heterogeneous resources, similar to stream processing, but also offers the elasticity and fault tolerance properties of batch processing. We present Ray Data, an implementation of the streaming batch model that improves throughput on heterogeneous batch inference pipelines by 3–8$\times$ compared to traditional batch and stream processing systems. When training Stable Diffusion, Ray Data matches the throughput of single-node ML data loaders while additionally leveraging distributed heterogeneous clusters to further improve training throughput by 31%.
11footnotetext: Work done while at Anyscale.
@@ 1 Introduction
Data processing is critical to machine learning applications. While model training and inference are both GPU-intensive, they also require significant I/O and CPU to load and preprocess datasets. CPU-based preprocessing is often the bottleneck in both training [30] and batch inference [22]. Meanwhile, as ML models evolve, the data processing functionality required has also become more diverse, spanning many modalities, transformations, and resource requirements [19, 41, 32].
There are numerous frameworks designed to scale out CPU-based data processing, typically based on either the batch [5, 47, 11] or stream [28, 9, 23, 2] processing models. These systems allow users to express a dataflow of logical operators (Figure 1), while the system automatically handles data distribution, task scheduling, and fault tolerance. Data processing in ML pipelines often consists of pure map transforms and thus can easily be expressed with this API [30]. However, these systems were designed for homogeneous CPU-only clusters. Heterogeneous ML pipelines have two key properties that limit performance and scalability when executed on CPU-centric data processing frameworks.
@@
(b) Training Stable Diffusion.
@@
Figure 1: Heterogeneous applications in batch inference and training, represented as logical dataflow graphs (nodes are operators). (a) represents a typical pipeline for video or image generation. (b) represents a pipeline for training stable diffusion model. In (b), UNet model is replicated for data-parallel training.
First, different operators require different degrees of physical parallelism, i.e. the number of concurrently executing instances. Thus, a key system requirement is to decouple the physical parallelism of each operator. For example, one may use many CPU threads to download an input dataset from cloud storage to match the throughput of a single GPU. This requires buffering intermediate data in memory until the downstream operator is ready to consume them. As data modalities such as text, images, and video proliferate [32, 40, 16, 38], the intermediate data size can be significant and unpredictable. Meanwhile, memory is more challenging to multiplex, as oversubscription is costly. Maintaining throughput while staying under memory limit is critical.
Second, some resources are more valuable than others. Thus, it is critical to isolate failure domains to the failed resource. For example, GPUs are more expensive than CPUs, but CPU failures are more likely because they are deployed in higher numbers and more often on spot instances [46]. Ideally, a CPU failure should have little impact on GPU execution. This is challenging to achieve simultaneously with exactly-once record processing and minimal run-time overheads. Even with stateless transforms, the system must track which records have been processed for each operator to avoid duplicating or dropping records during failover.
Thus, a distributed data processing system for ML pipelines must: (1) dynamically adjust each operator’s parallelism according to actual compute and memory usage, while (2) minimizing run-time and recovery overheads from failures and dynamic re-scaling. Current batch and stream processing systems can achieve one but not both. Batch processing systems execute operators in synchronous stages of stateless and immutable tasks, replaying them upon failure. This method, known as lineage reconstruction, offers high fault tolerance and elasticity but limited support for dynamic parallelism [11, 47]. Stream processing systems execute operators with asynchronous stateful executors that can decide locally when to materialize records. However, this decentralized approach requires a recovery method that imposes either high run-time overhead from logging [2, 28] or high recovery overhead from global checkpointing and rollback [9, 28].
To address these challenges, we present the streaming batch model, a hybrid model for efficient and fault-tolerant heterogeneous execution. Similar to the batch processing model, a centralized scheduler partitions data across a stage of stateless tasks. Tasks can run on any executor and are recovered via lineage reconstruction, enabling elasticity and fault tolerance while avoiding unnecessarily expensive logging. However, similar to the stream processing model, stages are executed asynchronously, allowing data to be streamed to the next operator and dynamically repartitioned.
(a) Batch processing model.
(b) Stream processing model.
(c) Streaming batch model (Ray Data).
@@
We present Ray Data, a distributed streaming batch system for heterogeneous workloads such as batch inference and ML training. Ray Data implements a logical dataflow API with an ahead-of-time execution planner and a run-time scheduler. The execution planner decides the initial number of partitions per operator. During execution, operators decide when to materialize a task’s intermediate partitions, allowing dynamic repartitioning based on local memory usage. Meanwhile, the centralized scheduler maintains a global view of running tasks and materialized partitions and can accordingly adjust the physical parallelism of each operator, i.e. its memory and compute allocation, while enforcing overall memory limits.
Ray Data uses Ray [27] as a distributed task backend. We extend Ray’s lineage-based recovery with a key feature needed for streaming batch execution: dynamic repartitioning.
We evaluate Ray Data on batch inference and ML training workloads that span diverse resource requirements (CPUs vs. GPUs), storage (local disk vs. cloud), and modalities (image vs. video). Ray Data outperforms batch and stream processing systems such as Spark and Flink with 3–8$\times$ better throughput. Ray Data is also able to match the throughput of single-node ML-specific data loaders such as tf.data and PyTorch DataLoader while additionally leveraging distributed and heterogeneous clusters. On a Stable Diffusion training benchmark, Ray Data can improve training time by 31% by leveraging a pool of 72 heterogeneous GPUs. In summary, we contribute:
  • The streaming batch model, an efficient and fault-tolerant execution model for distributed heterogeneous processing.
  • An online and heterogeneity-aware scheduling policy for streaming batch systems that can enforce total memory limits and maximize total compute utilization.
  • Ray Data, a fault-tolerant, memory-efficient, and autotuning implementation of the streaming batch model.
@@ 2 Background
We overview key aspects and limitations of the batch and stream processing models, including ML data loaders, by analyzing how effectively each system can:
  • Manage memory for intermediate data between operators.
  • Maximize utilization of heterogeneous compute resources.
@@ 2.1 Applications
We target ML training and inference pipelines that require data pre- and post-processing using CPUs, GPUs, or both. Our goal is maximizing throughput, while providing sub-second latency when used as data loaders for ML training [30]. Similar to current ML dataloaders [30, 35], we primarily target map-style per-row transforms. Operations that require all-to-all shuffle exchanges, such as sort and group-by, are also supported in the system but are already discussed in [25].
Figure 1(a) shows a typical batch inference pipeline, which uses CPUs to load and filter a dataset from cloud storage (e.g., HDFS or S3), GPUs to produce predictions, then CPUs to upload the results. A key characteristic is that the different operators may require different degrees of parallelism. For example, load may require many CPU threads to download an input dataset from cloud storage, while filter may only require one per core. Also, CPUs typically vastly outnumber GPUs in a cluster, so predict runs at much lower parallelism.
Figure 1(b) shows a typical distributed training pipeline for the Stable Diffusion model [38]. The pipeline uses CPUs to load and preprocess image-text pairs, GPUs to produce encodings with a pre-trained Encoder model, then GPUs to train a UNet model. Even GPU operators can leverage heterogeneity: colocating Encoder and UNet on the same GPUs can lead to lower training throughput, as it takes valuable resources from the UNet. If the operators’ physical resources are decoupled, then Encoder can be run on lower-end GPUs. This reduces overall cost by increasing utilization on the UNet GPUs.
@@ 2.2 Batch Processing Model
Batch processing systems are designed to allow a task to run and therefore recover on any executor. The system transforms the user-provided logical DAG (Figure 1(a)) into a physical DAG of stages of data-parallel tasks (Figure 2(a)). Tasks are stateless and materialize their input and output partition(s). This is key to lineage-based recovery, which logs only the logical DAG and re-executes tasks to recover lost partitions. Elastic scaling is supported by simply adding or removing executors. Examples include MapReduce [11], Apache Hadoop [5], Apache Spark [47], Spark Streaming [48], and Apache Flink in BATCH execution mode [9].
However, to make this recovery method practical, the system imposes two significant restrictions on execution. First, each stage must fully execute and materialize its outputs before executing the next stage. This simplifies scheduling and recovery, as a (re)scheduled task never idles waiting for its inputs. To reduce overheads from materialization, consecutive map operators are often fused into one stage [47]. This is effective when operators require the same resources. For example, suppose the load and filter operators in Figure 1(a) are fused. If each uses 1 CPU, then this has no impact on compute utilization, but greatly reduces the memory footprint, as data can be loaded and filtered one batch at a time.
Unfortunately, when different stages require different resources, operator fusion can cause significant underutilization. For example, with 4 CPUs and 2 GPUs, if Figure 1(a) fused all operators, then we could execute at most two tasks at a time. This would underutilize the CPUs while leaving the GPUs idle during CPU execution. Disabling fusion of heterogeneous operators, as in Figure 2(a), is also imperfect, as it requires materializing each stage’s outputs. This prevents pipelining between CPUs and GPUs, while producing high memory pressure and likely disk spilling.
Second, the data partitioning must be determined before execution so that it can be recorded in the lineage. This prevents the system from using run-time information such as the in-memory size of intermediate data rows when deciding the partitioning strategy. Thus, even if pipelined stage execution were supported, there can still be high memory pressure from individual intermediate partitions that are too large. For example, suppose that Figure 1(a) were run on a dataset of 100 videos, each 10MB on disk but 20GB decoded [7]. The system may choose 100 partitions to ensure good load-balancing for load, but each task would require 20GB memory! While some control is exposed to the user, typically one can only specify a target number [47], again before execution. Thus, it is left to the user to predict each operator’s memory usage, then manually configure the number of partitions.
@@ 2.3 Stream Processing Model
Stream processing architectures optimize for online scenarios and asynchronous execution. Examples include Naiad [28], Apache Flink [9], Spark Continuous Processing [42], MillWheel [2], and Apache Kafka [23]. Typically, each logical operator is assigned a parallelism, by the user or the system, which determines how many physical operator instances to create. For example, the parallelism of A-B-C in Figure 2(b) is 4-2-4. Consecutive logical operators that have the same resource requirements and parallelism are fused to avoid materialization. Physical operators execute asynchronously on their input stream(s) and exchange record batches directly, without involving a centralized scheduler.
Each physical operator processes a pre-determined stream partition. For example, in Figure 2(b), A8 is pre-assigned to CPU3. In contrast, Figure 2(a) executes A8 earlier, on any free CPU. However, unlike batch processing systems, physical operators are stateful and can choose at run time how many records to process at a time. For example, in Figure 2(b), the system assigns CPU3 1/4 of the total key range for operators A and C, but the executor chooses when to materialize A4 vs. A8 and C4 vs. C8. Typically, the executor will accumulate records up to a time or memory limit, then materialize and send the batch to a downstream executor. If the downstream executor is overloaded, backpressure is applied to limit memory.
Physical operators are tied to an executor and assigned one global partition, so reconfiguration is important for load-balancing. Reconfiguration efficiency is a common challenge [43], and especially so for heterogeneous pipelines. First, for heterogeneous operators, fusion is often impractical. For example, Figure 2(b) does not fuse A and C because B requires a GPU. Determining the optimal parallelism for A and C can be challenging without run-time information. Second, heterogeneous clusters can greatly reduce cost, but efficient elastic scaling and failure handling are critical.
There is a common tradeoff among stream processing systems between run-time overheads vs. reconfigurability and recovery overheads. Many systems use asynchronous global checkpointing [28, 9]. This typically imposes low execution overheads, but any failure causes a global rollback to the last checkpoint. Also, global checkpoints must coordinate each process’s local checkpoints, so static membership is often assumed for simplicity [13]. Thus, reconfiguration typically introduces long pauses, as it requires taking then restarting from a global checkpoint with the new configuration [9].
Other systems use logging [2, 23]. This allows for fast recovery via log replay and fast repartitioning of physical operators. However, intermediate records must be durably logged before releasing to the downstream operator, adding higher run-time overheads. This is acceptable for online systems that interact frequently with the external world, but ML pipelines are typically offline systems where most intermediate values are safe to rollback and thus do not need durability.
@@
ML-specific data loaders are single-node systems that maximize I/O and CPU bandwidth to improve local GPU utilization. Examples include tf.data [30] and PyTorch DataLoader [35]. These systems can be viewed as a special case of stream processing: they launch a fixed pool of worker threads or processes at run time, which continuously load, preprocess, and feed data to an end GPU consumer. The pool must be tuned to match the GPU’s throughput without running out of memory. tf.data automatically adjusts operator parallelism at run time to maximize GPU utilization [30].
However, in general these data loaders are narrow in scope: (1) they do not support distributed execution, and (2) GPUs are always assumed to be sinks. (1) prevents load-balancing and heterogeneous clusters. Both (1) and (2) make it impossible to leverage heterogeneous GPUs in the Stable Diffusion example in Figure 1(b), as well as batch inference pipelines that alternate CPU and GPU operators. Adapting data loaders to support these applications would require effort equivalent to re-building a distributed data processing framework.
Batch [11, 5, 47] Stream [28, 9, 2, 23] PyTorch DL [35] tf.data [30] Streaming batch (Ray Data)
Automatic partitioning $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Dynamic repartitioning $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$: streaming repartition
Min. pipeline granularity Stage Partition Partition Partition Partition
Dynamic parallelism $\checkmark$ $\checkmark$/$\times$ $\times$ $\checkmark$ $\checkmark$
Distributed execution $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@
Min. rollback granularity Partition Record/Epoch Epoch Epoch Partition
Table 1: Features for heterogeneous and distributed processing. Dynamic repartitioning is important for memory efficiency. Finer pipeline granularity improves utilization for heterogeneous resources. Dynamic parallelism is important for adaptivity. Stream processing systems use logging for dynamic parallelism with zero downtime and record-level rollback, or checkpointing, which requires a checkpoint to reconfigure and global rollback on failure (§ 2.3).

### v5

@@ Abstract
While ML model training and inference are both GPU-intensive, CPU-based data processing is often the bottleneck. Distributed data processing systems based on the batch or stream processing models assume homogeneous resource requirements. They excel at CPU-based computation but either under-utilize heterogeneous resources or impose high overheads on failure and reconfiguration.
[…]
footnotetext: Equal contribution.
@@ 1 Introduction
Data processing is critical to machine learning (ML) applications. While ML workloads are known to be GPU-intensive, they also require significant I/O and CPU to load, preprocess, and move data to the GPU. Indeed, CPU-based preprocessing is often the bottleneck in both training [28] and batch inference [20]. Meanwhile, as ML models evolve, the data processing functionality required has also become more diverse, spanning many data modalities such as text, image, and video. These new workloads introduce complex transformations and resource requirements [16, 38, 30]. As ML models have grown larger in scale, there is also a critical need to scale the data processing alongside the ML workload.
@@
(b) Multimodal training with Stable Diffusion [35].
@@
Figure 1: Logical dataflow graphs representing heterogeneous ML applications. Nodes are operators. (a) A pipeline for video or image generation. (b) A multimodal training pipeline. The UNet model is replicated for distributed data-parallel training.
At first glance, such applications seem simple to scale: both distributed training and batch inference can be expressed as dataflow graphs (Figure˜1) where each node is an operator that executes a transform over its incoming data. The data transforms used are often embarrassingly parallel, e.g., randomly cropping one image for training. There have been numerous frameworks built to address this exact problem of scaling arbitrary distributed dataflows [8, 4, 43, 17, 27, 7, 21, 1]. Unfortunately, none of these are able to fully address the problem of efficient scaling of heterogeneous dataflows that use a mix of resources, e.g., GPUs and CPUs.
Resource heterogeneity requires pipelining execution across different operators to maximize overall utilization. For example, maximizing GPU utilization requires pipelining GPU compute with any CPU and I/O operations needed for data preprocessing, postprocessing, and loading. This introduces two requirements. First, the system must manage memory for intermediate data between operators that require different resources. This is challenging because the system must maximize overall throughput while ensuring sufficient overall memory. If too little intermediate data is buffered, the GPU will idle, wasting valuable resources. On the other hand, if too much data is buffered, the system may run out of memory or incur expensive overheads from swapping intermediate data to disk. This is especially important in multimodal applications, which often produce large intermediate outputs, e.g., one video file produces a series of decoded frames.
Second, achieving efficiency in a heterogeneous setting requires elasticity to keep the overall pipeline balanced, as the optimal resource allocation per data record and operator may not be known until run time. For example, in video generation, longer videos will take longer to process, so simple schemes like round-robin assignment of videos to resources can result in poor utilization. Ideally, the system should also be able to dynamically reallocate each operator’s resources based on actual load. Similarly, in the distributed setting, the system should be able to add and remove physical resources to the cluster with minimal downtime, including in unexpected cases such as node failures.
While there have been numerous frameworks built to scale traditional CPU-based data analytics, none achieve both of these requirements. Typically, they fall into one of two categories: batch or streaming systems. Batch systems [8, 43, 4, 17] divide the dataset into partitions, i.e. batches of data records (Figure˜2(a)). Partitions can be processed by any executor, enabling elasticity and fine-grained recovery. However, batch systems process operators one at a time, which prevents pipelining and adds high memory overhead (Figure˜3a). Streaming systems [27, 7, 21, 1] make the opposite tradeoff: they allow executors to buffer and send data directly to each other in dynamically sized batches (Figure˜2(b)). This results in better utilization for heterogeneous resources and lower memory usage (Figure˜3b). However, each operator and data range is bound to a specific set of resources, making it difficult to dynamically load-balance inputs, adjust resource allocations per operator, and scale the cluster.
We present Ray Data, a streaming batch system for heterogeneous workloads spanning batch inference and ML training. The key idea is to use partitions as the unit of execution to achieve elasticity, but to allow partitions to be dynamically created and streamed between heterogeneous operators for memory-efficient pipelining (Figure˜2(c)). In Ray Data, a partition is a dynamically sized batch of records. Each partition is consumed by one or more tasks, each of which executes one (possibly fused) operator and dynamically produces one or more output partitions. Tasks are dynamically assigned to resources at run time, enabling elasticity as in batch systems while supporting pipelined execution across heterogeneous resources as in streaming systems (Figure˜3c).
(a) Batch systems.
(b) Streaming systems.
(c) Streaming batch systems.
@@
The key challenges stem from reconciliation of batch and streaming execution. First, batch systems also use partitions for resource elasticity but the partitions are typically fixed before run time. This is so that the system can correctly recover the partition in case of failure. However, because partition size determines memory footprint, this can lead to unpredictable memory usage at run time. Ray Data must therefore size each partition based on real-time memory usage. Second, when multiple operators require the same physical resource, avoiding stalls and data spills requires joint control of memory allocation and compute resources.
To address these challenges, Ray Data uses a centralized scheduler that dynamically partitions the dataset and assigns tasks to consume partitions. Executors can further dynamically partition their outputs based on actual memory usage to enforce cluster memory limits. The Ray Data scheduler uses an online and memory-aware task scheduling policy, enabling joint control over cluster memory and compute resources at partition granularity. To maintain scalability, Ray Data builds on Ray [26] to decouple the control and data planes: the Ray Data scheduler issues partition control decisions but the partition data never flows through it.
[…]
  • The streaming batch model for memory-efficient and resource-elastic execution of heterogeneous distributed dataflows.
  • An online scheduling policy for heterogeneous data processing that maximizes total compute utilization while enforcing total memory limits.
  • Ray Data, an efficient, scalable, and fault-tolerant streaming batch system for heterogeneous ML applications.
Figure 3: Execution timelines for Figure˜2. (a) Batch systems dynamically assign partitions to resources but execute one stage at a time, materializing all intermediate partitions. (b) Streaming systems pipeline records between executors in dynamically sized batches but statically assign resources and records to executors, which can reduce utilization. (c) Streaming batch systems dynamically size partitions and dynamically assign partitions to resources for efficient pipelining and load-balancing.
@@ 2 Background
We overview key aspects and limitations of the batch and stream processing models (Table˜1), including ML data loaders, by analyzing how effectively each system can:
  • Maximize utilization of heterogeneous compute.
  • Manage memory for intermediate data.
@@ 2.1 Applications
We target throughput for ML training and inference pipelines that require data pre- and post-processing using CPUs, GPUs, or both. Similar to current ML dataloaders [28, 33], we primarily target map-style per-row transforms. We support operations that require all-to-all shuffle exchanges, such as sort and group-by, through techniques described in [24].
Figure˜1(a) shows a typical batch inference pipeline, which uses CPUs to load a dataset from cloud storage, GPUs to produce predictions, then CPUs to upload the results. Figure˜1(b) shows a typical multimodal distributed training pipeline, in this case for Stable Diffusion [35]. The pipeline uses CPUs to load image-text pairs, GPUs to run inference on frozen Encoder models, and GPUs to train a UNet model. Note that GPU operators can also benefit from heterogeneity and disaggregation: placing the Encoders on cheaper GPUs can reduce cost by allowing UNet more GPU memory (Section˜5.2.2).
@@ 2.2 Batch Processing Model
Batch processing systems use stateless tasks to allow a task to run and therefore recover on any executor. Examples include MapReduce [8], Apache Hadoop [4], Apache Spark [43], Spark Streaming [44], and Apache Flink in BATCH execution mode [7]. Before execution, the system transforms the logical DAG (Figure˜1(a)) into a DAG of tasks and data partitions (Figure˜2(a)). Tasks are stateless and materialize their input and output partition(s). This is key to lineage-based recovery, which avoids logging data by recording only the DAG and re-executing tasks to recover lost partitions. Elastic scaling is supported by simply adding or removing executors.
However, to make this recovery method practical, the system imposes two significant restrictions on execution. First, each stage must fully execute before executing the next stage (Figure˜3a). This simplifies scheduling and recovery, as a (re)scheduled task never idles waiting for its inputs, but prevents pipelining and requires materializing all outputs between stages. The consequences are not as severe in homogeneous applications because consecutive map operators that require the same resources can be fused into one stage (Figure˜2(a)). In the heterogeneous setting, fusion of operators requiring different resources couples their parallelism, easily leading to under-utilization in heterogeneous clusters.
[…]
@@ 2.3 Stream Processing Model
Stream processing architectures optimize for low latency by using stateful executors that exchange records directly without involving a centralized scheduler. Examples include Naiad [27], Apache Flink [7], Spark Continuous Processing [39], MillWheel [1], and Apache Kafka [21]. Before execution, the system assigns resources to each executor, and then shards each logical operator across the executors. For example, in Figure˜2(b) CPU0 is assigned the even indices for all CPU operators. Executors execute asynchronously: each can buffer records and decide when to send a batch to a downstream operator. If the downstream executor is overloaded, backpressure is applied to limit overall memory usage. This enables memory-efficient pipelining across heterogeneous operators.
However, this is achieved by coupling each executor to a set of resources, operator(s), and data range(s), which prevents dynamic load balancing. Thus, cluster reconfiguration requires expensive protocols to maintain distributed state consistency. Some systems use global checkpointing [27, 7], which minimizes run-time overheads. However, any failure or reconfiguration event triggers a global rollback to the last checkpoint [11, 7]. Other systems use logging [1, 21], which enables fast recovery and reconfiguration by durably logging intermediate data. However, this adds high run-time overheads [27] that are unacceptable for ML pipelines, where most computations are idempotent and thus do not need durability.
@@
ML data loaders such as tf.data [28] and PyTorch DataLoader [33] use CPUs to load data onto a local GPU. They can be viewed as single-node stream processing systems: they launch a fixed pool of worker threads or processes that continuously load, preprocess, and feed data to a GPU consumer. ML data loaders do not typically support distributed execution, which prevents load-balancing across nodes and the use of heterogeneous node types. Also, GPUs are assumed to be sinks, which precludes use cases such as batch inference that require CPU-based post-processing in addition to preprocessing.
Batch Stream Streaming batch
[8, 4, 43, 17] [27, 7, 1, 21] PyTorch DL [33] tf.data [28] (Ray Data)
Dynamic partitioning $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Dynamic resource assignment $\checkmark$ $\times$ $\times$ $\checkmark$ (local only) $\checkmark$
@@
Min. rollback granularity Partition Record/Epoch Job Epoch Partition
Heterogeneous node types $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
Table 1: Systems for heterogeneous data processing. Dynamic partitioning: the system decides how to batch records between operators. Dynamic resource assignment: the system dynamically assigns cluster resources to operators. Rollback granularity: the amount that needs to be re-executed upon failure or cluster reconfiguration. Streaming systems use logging for record-level rollback [1, 21] or global checkpointing for epoch-level rollback [27, 7]. Some ML data loaders support dynamic partitioning and resource assignment [28], but only within a node.